An organic acid is composed of carbon (62.04%), hydrogen (10.43%), and oxygen (27.55%). Its molar mass is 116.16 g/mol. Determine the molecular formula of the compound.
Answer #1

1)

we have mass of each elements as:

C: 62.04 g

H: 10.43 g

O: 27.55 g

Divide by molar mass to get number of moles of each:

C: 62.04/12.01 = 5.1657

H: 10.43/1.008 = 10.3472

O: 27.55/16.0 = 1.7219

Divide by smallest to get simplest whole number ratio:

C: 5.1657/1.7219 = 3

H: 10.3472/1.7219 = 6

O: 1.7219/1.7219 = 1

  

So empirical formula is:C3H6O

Molar mass of C3H6O,

MM = 3*MM(C) + 6*MM(H) + 1*MM(O)

= 3*12.01 + 6*1.008 + 1*16.0

= 58.078 g/mol

Now we have:

Molar mass = 116.16 g/mol

Empirical formula mass = 58.078 g/mol

Multiplying factor = molar mass / empirical formula mass

= 116.16/58.078

= 2

So molecular formula is:C6H12O2

Answer: C6H12O2